OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (anonymous):

First find f(g(x))

OpenStudy (anonymous):

Do you know how to find f(g(x))?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

when i set it up its weird

OpenStudy (anonymous):

can you show me please

OpenStudy (anonymous):

Yeah sure, \[f(g(x)) = (x^2-6x-7)+8\]

OpenStudy (anonymous):

Just sub in 2 now where ever the x is and you'll have f(g(2)).

OpenStudy (anonymous):



OpenStudy (anonymous):

got it its -7

OpenStudy (anonymous):

Yeah that's right.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). first find f(g(x)) how you do this is by putting g(x) value into f(x) as in.. f(g(x)) = x + 8 where x -> x^2 - 6x -7 = x^2 - 6x - 7 + 8 now replace the value of x by 2(because f(g(2)) 4 - 12 - 7 + 8 = -7 :) Oh look.. I'm too late to answer.. oh well.